x-2 over 5 = 3 over 10
Answer by stanbon(75887)

(x-2)/5 = 3/10
Cross-multiply to get:
10(x-2) = 15
10x-20 = 15
10x = -5
x = -1/2
==============
Cheers,
Stan H.